# ObjectCompose: Evaluating Resilience of Vision-Based Models on Object-to-Background Compositional Changes

### v2

@@ 0.A.14. Misclassified Samples
0.A.15. Limitations and Future Directions
0.A.16. Potential External Factors
0.A.17. Dataset Distribution
0.A.18. Evaluation on Background/Foreground Images
@@ 0.A.1 List of prompts
Class label |
"This is a picture of a class name" |
BLIP-2 Caption |
Captions generated from BLIP-2 image to caption |
Colorprompt-1 |
"This is a picture of a vivid red background" |
Colorprompt-2 |
"This is a picture of a vivid green background" |
Colorprompt-3 |
"This is a picture of a vivid blue background" |
Colorprompt-4 |
"This is a picture of a vivid colorful background" |
Textureprompt-1 |
"This is a picture of textures in the background" |
Textureprompt-2 |
"This is a picture of intricately textured background" |
Textureprompt-3 |
"This is a picture of colorful textured background" |
Textureprompt-4 |
"This is a photo of distorted textures in the background" |
Adversarial |
Captions generated from BLIP-2 image to caption. |
@@ 0.A.5 Ablation on Background Changes
In this section, we report results on ImageNet-B and COCO-DC for uni-modal classifiers in Table 8 and Table 9 reports zero-shot classification results on ImageNet-B. Furthermore, ablations across diverse color and texture prompts is provided in Table 10 and 11.
@@ 0.A.6 Evaluation on Adversarially Trained models
In this section, we evaluate adversarially trained Res-18, Res-50, and Wide-Res-50 models across background changes induced by our methods and baseline methods (See Table 12, 13, and 14).
| Datasets | ||||||||
| =0.5 | =2.0 | =4.0 | =8.0 | =0.5 | =2.0 | =4.0 | =8.0 | |
| Original | 88.00 | 78.30 | 69.70 | 54.40 | 84.60 | 81.40 | 68.80 | 57.50 |
| ImageNet-E (=-20) | 84.50 | 77.01 | 69.10 | 54.80 | 83.40 | 80.00 | 68.80 | 59.00 |
| ImageNet-E (=20) | 81.41 | 74.94 | 66.36 | 52.52 | 81.61 | 75.75 | 65.65 | 55.05 |
| ImageNet-E () | 75.15 | 66.16 | 56.36 | 45.35 | 72.82 | 66.66 | 55.75 | 45.05 |
| LANCE | 76.37 | 66.78 | 59.13 | 45.17 | 74.99 | 73.19 | 61.60 | 48.68 |
| Class label | 87.10 | 79.90 | 69.40 | 57.30 | 85.00 | 79.90 | 70.90 | 57.80 |
| BLIP-2 Caption | 80.90 | 73.10 | 67.10 | 51.00 | 79.50 | 75.30 | 63.10 | 53.40 |
| Color | 56.90 | 45.80 | 35.40 | 25.70 | 53.20 | 46.80 | 32.80 | 23.40 |
| Texture | 59.20 | 47.10 | 38.60 | 28.70 | 54.30 | 48.10 | 35.50 | 26.20 |
| Adversarial | 12.10 | 19.80 | 24.60 | 26.80 | 10.90 | 12.40 | 16.90 | 17.20 |
| Datasets | ||||||||
| =0.5 | =2.0 | =4.0 | =8.0 | =0.5 | =2.0 | =4.0 | =8.0 | |
| Original | 95.20 | 89.30 | 83.20 | 72.40 | 94.30 | 91.10 | 80.90 | 70.80 |
| ImageNet-E (=-20) | 93.10 | 89.20 | 82.00 | 70.70 | 91.70 | 88.50 | 79.60 | 69.10 |
| ImageNet-E (=20) | 92.52 | 86.36 | 80.60 | 67.97 | 90.40 | 86.96 | 76.26 | 68.08 |
| ImageNet-E () | 84.44 | 78.78 | 71.71 | 58.58 | 80.50 | 76.76 | 65.75 | 56.86 |
| LANCE | 81.94 | 78.96 | 70.11 | 59.72 | 83.52 | 80.46 | 69.83 | 61.32 |
| Class label | 92.40 | 88.50 | 82.70 | 72.50 | 90.70 | 88.60 | 80.20 | 70.50 |
| BLIP-2 Caption | 87.90 | 83.70 | 79.00 | 67.90 | 86.60 | 84.60 | 73.70 | 65.70 |
| Color | 70.80 | 60.30 | 53.20 | 39.50 | 67.20 | 58.50 | 44.40 | 34.20 |
| Texture | 69.70 | 61.00 | 54.60 | 43.40 | 64.90 | 59.70 | 48.00 | 37.70 |
| Adversarial | 10.80 | 17.10 | 18.10 | 16.60 | 10.70 | 11.90 | 14.70 | 13.40 |
| Datasets | ||||||||
| =0.5 | =2.0 | =4.0 | =8.0 | =0.5 | =2.0 | =4.0 | =8.0 | |
| Original | 96.20 | 92.60 | 89.10 | 78.70 | 95.60 | 94.00 | 87.00 | 78.30 |
| ImageNet-E (=-20) | 94.10 | 91.10 | 86.60 | 76.40 | 93.20 | 91.60 | 84.20 | 76.90 |
| ImageNet-E (=20) | 92.82 | 89.29 | 83.53 | 75.15 | 91.21 | 88.68 | 81.11 | 73.63 |
| ImageNet-E () | 87.17 | 76.56 | 66.16 | 82.92 | 82.22 | 82.00 | 71.41 | 62.42 |
| LANCE | 84.18 | 81.34 | 77.07 | 64.51 | 83.48 | 83.09 | 77.40 | 66.09 |
| Class label | 93.50 | 90.60 | 87.30 | 78.80 | 92.10 | 90.80 | 83.70 | 75.50 |
| BLIP-2 Caption | 90.20 | 86.20 | 82.80 | 74.30 | 88.90 | 86.70 | 80.00 | 69.20 |
| Color | 72.2 | 66.60 | 60.70 | 51.20 | 68.10 | 65.10 | 51.40 | 40.90 |
| Texture | 73.80 | 66.70 | 61.70 | 53.60 | 67.20 | 64.60 | 52.00 | 44.20 |
| Adversarial | 13.90 | 22.90 | 28.00 | 32.00 | 12.40 | 15.10 | 19.50 | 20.60 |
@@ 0.A.14 Misclassified Samples
0.A.15 Limitations and Future Directions
Limitations: In Figure 33, we observe that for objects covering a small region in the image, relying solely on the class name to guide the diffusion model can result in alterations of the object shape, expanding the influence of the class name semantics to larger image regions. However, by supplementing with descriptive captions that encompass the object-to-background context, we partially mitigate this effect. Furthermore, the generated textured background can inadvertently camouflage the object. To address this concern, we slightly expand the object mask to clearly delineate the object boundaries.
Future Directions: Our current work represents one of the preliminary efforts in utilizing diffusion models to study the object-to-background context in vision-based systems. Based on our observations and analysis, the following are the interesting future directions.
0.A.16 Potential External Factors
@@
Extension of Objects: As already detailed in Appendix 0.A.15, we encountered challenges when dealing with objects that occupy a small region in the image, sometimes leading to certain unwanted extensions to objects. To mitigate this, we filtered our dataset to focus on images where the object covers a significant area. Additionally, we slightly expand object masks computed using SAM to better define boundaries and prevent object shape distortion in the background.
@@ 0.A.17 Dataset Distribution
0.A.18 Evaluation on Background/Foreground Images

### v4

@@ 0.A.14. Misclassified Samples
0.A.15. Potential External Factors
0.A.16. Dataset Distribution and Comparison
0.A.17. Evaluation on Background/Foreground Images
0.A.18. Insights
0.A.19. Calibration Metrics
0.A.20. Ablation on Adversarial Loss
0.A.21. Reproducibility and Ethics Statement
@@ 0.A.1 List of prompts
| Class label | "This is a picture of a class name" |
| BLIP-2 Caption | Captions generated from BLIP-2 image to caption |
| Colorprompt-1 | "This is a picture of a vivid red background" |
| Colorprompt-2 | "This is a picture of a vivid green background" |
| Colorprompt-3 | "This is a picture of a vivid blue background" |
| Colorprompt-4 | "This is a picture of a vivid colorful background" |
| Textureprompt-1 | "This is a picture of textures in the background" |
| Textureprompt-2 | "This is a picture of intricately textured background" |
| Textureprompt-3 | "This is a picture of colorful textured background" |
| Textureprompt-4 | "This is a photo of distorted textures in the background" |
| Adversarial | Captions generated from BLIP-2 image to caption. |
@@ 0.A.3 Object Distortion in LANCE
We calculate the FID score by comparing the background changes applied on our ImageNet-B dataset with the original ImageNet val. set (Tab. 8). Our background modifications such as Class Label, BLIP Caption, & Color achieve FID scores close to the original images, while our more complex background changes (Texture, Adversarial) show significant improvement over related works[44, 61].
@@ 0.A.5 Ablation on Background Changes
In this section, we report results on ImageNet-B and COCO-DC for uni-modal classifiers in Table 9 and Table 10 reports zero-shot classification results on ImageNet-B. Furthermore, ablations across diverse color and texture prompts is provided in Table 11 and 12.
@@ 0.A.6 Evaluation on Adversarially Trained models
In this section, we evaluate adversarially trained Res-18, Res-50, and Wide-Res-50 models across background changes induced by our methods and baseline methods (See Table 13, 14, and 15).
| Datasets | |||||||||
| =0.5 | =2.0 | =4.0 | =8.0 | =0.5 | =2.0 | =4.0 | =8.0 | ||
| Original | 88.00 | 78.30 | 69.70 | 54.40 | 84.60 | 81.40 | 68.80 | 57.50 | |
| ImageNet-E (=-20) | 84.50 | 77.01 | 69.10 | 54.80 | 83.40 | 80.00 | 68.80 | 59.00 | |
| ImageNet-E (=20) | 81.41 | 74.94 | 66.36 | 52.52 | 81.61 | 75.75 | 65.65 | 55.05 | |
| ImageNet-E () | 75.15 | 66.16 | 56.36 | 45.35 | 72.82 | 66.66 | 55.75 | 45.05 | |
| LANCE | 76.37 | 66.78 | 59.13 | 45.17 | 74.99 | 73.19 | 61.60 | 48.68 | |
| Class label | 87.10 | 79.90 | 69.40 | 57.30 | 85.00 | 79.90 | 70.90 | 57.80 | |
| BLIP-2 Caption | 80.90 | 73.10 | 67.10 | 51.00 | 79.50 | 75.30 | 63.10 | 53.40 | |
| Color | 56.90 | 45.80 | 35.40 | 25.70 | 53.20 | 46.80 | 32.80 | 23.40 | |
| Texture | 59.20 | 47.10 | 38.60 | 28.70 | 54.30 | 48.10 | 35.50 | 26.20 | |
| Adversarial | 12.10 | 19.80 | 24.60 | 26.80 | 10.90 | 12.40 | 16.90 | 17.20 | |
| Datasets | |||||||||
| =0.5 | =2.0 | =4.0 | =8.0 | =0.5 | =2.0 | =4.0 | =8.0 | ||
| Original | 95.20 | 89.30 | 83.20 | 72.40 | 94.30 | 91.10 | 80.90 | 70.80 | |
| ImageNet-E (=-20) | 93.10 | 89.20 | 82.00 | 70.70 | 91.70 | 88.50 | 79.60 | 69.10 | |
| ImageNet-E (=20) | 92.52 | 86.36 | 80.60 | 67.97 | 90.40 | 86.96 | 76.26 | 68.08 | |
| ImageNet-E () | 84.44 | 78.78 | 71.71 | 58.58 | 80.50 | 76.76 | 65.75 | 56.86 | |
| LANCE | 81.94 | 78.96 | 70.11 | 59.72 | 83.52 | 80.46 | 69.83 | 61.32 | |
| Class label | 92.40 | 88.50 | 82.70 | 72.50 | 90.70 | 88.60 | 80.20 | 70.50 | |
| BLIP-2 Caption | 87.90 | 83.70 | 79.00 | 67.90 | 86.60 | 84.60 | 73.70 | 65.70 | |
| Color | 70.80 | 60.30 | 53.20 | 39.50 | 67.20 | 58.50 | 44.40 | 34.20 | |
| Texture | 69.70 | 61.00 | 54.60 | 43.40 | 64.90 | 59.70 | 48.00 | 37.70 | |
| Adversarial | 10.80 | 17.10 | 18.10 | 16.60 | 10.70 | 11.90 | 14.70 | 13.40 | |
| Datasets | |||||||||
| =0.5 | =2.0 | =4.0 | =8.0 | =0.5 | =2.0 | =4.0 | =8.0 | ||
| Original | 96.20 | 92.60 | 89.10 | 78.70 | 95.60 | 94.00 | 87.00 | 78.30 | |
| ImageNet-E (=-20) | 94.10 | 91.10 | 86.60 | 76.40 | 93.20 | 91.60 | 84.20 | 76.90 | |
| ImageNet-E (=20) | 92.82 | 89.29 | 83.53 | 75.15 | 91.21 | 88.68 | 81.11 | 73.63 | |
| ImageNet-E () | 87.17 | 76.56 | 66.16 | 82.92 | 82.22 | 82.00 | 71.41 | 62.42 | |
| LANCE | 84.18 | 81.34 | 77.07 | 64.51 | 83.48 | 83.09 | 77.40 | 66.09 | |
| Class label | 93.50 | 90.60 | 87.30 | 78.80 | 92.10 | 90.80 | 83.70 | 75.50 | |
| BLIP-2 Caption | 90.20 | 86.20 | 82.80 | 74.30 | 88.90 | 86.70 | 80.00 | 69.20 | |
| Color | 72.2 | 66.60 | 60.70 | 51.20 | 68.10 | 65.10 | 51.40 | 40.90 | |
| Texture | 73.80 | 66.70 | 61.70 | 53.60 | 67.20 | 64.60 | 52.00 | 44.20 | |
| Adversarial | 13.90 | 22.90 | 28.00 | 32.00 | 12.40 | 15.10 | 19.50 | 20.60 | |
@@ 0.A.14 Misclassified Samples
0.A.15 Potential External Factors
@@
Extension of Objects: As already detailed in Appendix 0.A.18.1, we encountered challenges when dealing with objects that occupy a small region in the image, sometimes leading to certain unwanted extensions to objects. To mitigate this, we filtered our dataset to focus on images where the object covers a significant area. Additionally, we slightly expand object masks computed using SAM to better define boundaries and prevent object shape distortion in the background.
@@ 0.A.16 Dataset Distribution and Comparison
As shown in Tab. 20, our curated ImageNet-B dataset is the largest in terms of both the number of images and classes compared to closely related works [44, 34, 61]. In contrast to [44, 34, 61], we extend our analysis to object detection by introducing the COCO-DC dataset. Our proposed background changes on ImageNet-B & COCO-DC, enable us to evaluate on more than 70k samples for classification & 5k samples for detection. Our automated framework of delineating between foreground & background facilitates future dataset expansion.
0.A.17 Evaluation on Background/Foreground Images
@@
0.A.18 Insights
Across Architectures. Our results (Tab. 1,2, 5, & 16) show that CNNs perform better than transformers across various background changes. We note that the mixing of background & foreground token features through the global attention mechanism may result in the reliance of transformer models prediction on outlier/background tokens. This is validated when we evaluate transformers which are trained to prioritize learning more salient features[9] (Tab. 17 & 22), resulting in improved performance under background changes.
Across Training Methods. Analysis of adversarially trained models (Fig. 6 & Tab. 13, 14, 15) reveal their robustness is confined to adversarial background changes, leaving them vulnerable to other background variations. Similar behavior is observed for models trained on stylized ImageNet dataset(Tab. 5). However, self-supervised training of uni-modal models on large extensively curated datasets shows performance gain across background changes(Fig. 7). Similarly, for multi-modal models, we observe that stabilizing training on large-scale datasets (as in EVA-CLIP) leads to significant improvement in zero-shot performance across all background changes(Tab. 2,10, & 11).
Across Vision Tasks. The absence of object-to-background context during classification model training creates a significant vulnerability to background changes. In contrast, object detection & segmentation models (Tab. 5, Fig. 8 & 14), which explicitly incorporate object-to-background context during training, show notably better resilience to background variations. Based on the above insights, we discuss current limitations and future directions next.
0.A.18.1 Limitations.
In Figure 34, we observe that for objects covering a small region in the image, relying solely on the class name to guide the diffusion model can result in alterations of the object shape, expanding the influence of the class name semantics to larger image regions. However, by supplementing with descriptive captions that encompass the object-to-background context, we partially mitigate this effect. Furthermore, the generated textured background can inadvertently camouflage the object. To address this concern, we slightly expand the object mask to clearly delineate the object boundaries.
0.A.18.2 Future Directions.
Our current work represents one of the preliminary efforts in utilizing diffusion models to study the object-to-background context in vision-based systems. Based on our observations and analysis, the following are the interesting future directions.
0.A.19 Calibration Metrics
Model calibration refers to how well a model’s predicted confidence levels correspond to its actual accuracy. Confidence represents the probability a model assigns to its predictions, while accuracy measures how often those predictions are correct. For example, if a model predicts with 70% confidence, a well-calibrated model should have an actual accuracy close to 70%. To quantify this, we use the Expected Calibration Error (ECE). ECE works by dividing the predictions into bins based on confidence intervals (e.g., 60%-70%, 70%-80%). Within each bin, the average confidence and accuracy are calculated, and the ECE is the weighted average of the differences between these values. To visually evaluate model calibration, reliability diagrams are used. These diagrams plot predicted confidence against actual accuracy, allowing us to compare different models. A well-calibrated model will show points that lie close to the diagonal on these plots. In Figures 35 and 36, we plot the reliability diagrams for different convolutional and transformer-based models, respectively.
0.A.20 Ablation on Adversarial Loss
Tab.23 presents an ablation study on adversarial loss, showing optimizing both vision & text based emebedding, leads to the most performance drop.
| Text | 28.4 | 46.5 | 40.9 | 47.2 | 19.0 | 46.5 | 31.5 | 37.1 |
| Latent | 32.3 | 47.5 | 42.8 | 50.2 | 1.6 | 44.5 | 26.4 | 35.1 |
| Combined | 18.4 | 32.1 | 25.0 | 31.7 | 2.0 | 28 | 14.4 | 21.6 |
0.A.21 Reproducibility and Ethics Statement
Reproducibility Statement: Our method uses already available pre-trained models and the codebase is based on several open source implementations. We highlight the main components used in our framework for reproducing the results presented in our paper, a) Diffusion Inpainting Implementation: We use the open-source implementation of Stable-Diffusion-Inpainting method (https://github.com/huggingface/diffusers/blob/main/src/diffusers/) with available pretrained weights (Stable-Diffusion-v-1-2) for background generation. b) Image-to-Segment Implementation: We use the official open-source implementation of FastSAM (https://github.com/CASIA-IVA-Lab/FastSAM) to get the segmentation masks of filtered ImageNet dataset. c) Image-to-Text Implementation: We use the official open-source implementation of BLIP-2(https://github.com/salesforce/LAVIS/tree/main/projects/blip2) to get the captions for each image. We will also provide captions for each image in our dataset. d) Adversarial Attack: We intent to open-source our codebase and release the script for crafting adversarial examples. e) Dataset: In the paper, we describe the procedure of filtering the images from ImageNet and COCO val. set. Furthermore, we will provide the filtered datasets, object masks as well as prompts used to generate different backgrounds.
Ethics Statement: Our work focuses on evaluating resilience of current vision and language models against natural and adversarial background changes in real images. This work can be utilized by an attacker to generate malicious backgrounds on real images as well as generate adversarial backgrounds which can fool the deployed computer-vision systems. Nevertheless, we believe that our research will pave the way for improved evaluation protocols to assess the resilience of existing models. This, in turn, is likely to drive the development of enhanced techniques for bolstering the resilience of deployed systems. Since we are benchmarking vision and vision-language models using a subset of images from publicly available ImageNet and COCO datasets, it’s relevant to mention that these datasets are known to have images of people which poses a privacy risk and further it is known to have biases which can encourage social stereotypes. In the future, we intend to benchmark our models on a less biased dataset to mitigate these concerns and ensure a fair evaluation.